How do i solve 10x-5y-10x-3y i know i have to group the xs and the ys so that gives me 10x-10x-5y-3y whats next?
Answer by stanbon(75887)

How do i solve 10x-5y-10x-3y i know i have to group the xs and the ys so that gives me 10x-10x-5y-3y whats next?
10x-10x = 0
-5y-3y = -8y
------------------
Ans: -8y
====================================
Cheers,
Stan H.